# Existence of variational solutions to doubly nonlinear systems in nondecreasing domains

### v1

@@ 3.1. Technical lemmas
(3.5) $\left\{\begin{array}[]{ll}\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\to\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}&\text{strongly in }L^{1}(\Omega_{T},\mathds{R}^{N}),\\[5.0pt] u_{\ell}\to u&\text{a.e.~{}in }\Omega_{T},\end{array}\right.$
@@ 3.2. Mollification in time
We will use the following regularization with respect to time introduced by Landes [39]. Given $h>0$ and a function $v$, the mollification $[v]_{h}$ is constructed to satisfy the ordinary differential equation
@@
with initial condition $[v]_{h}(0)=v_{o}$. More precisely, let $X$ be a separable Banach space, $v_{o}\in X$ and $v\in L^{r}(0,T;X)$ for some $1\leq r\leq\infty$. Later on, we will mainly consider $X=L^{q}(\Omega,\mathds{R}^{N})$ and $X=W^{1,p}_{u_{\ast}}(\Omega,\mathds{R}^{N})$. For $h\in(0,T]$, we define the mollification
(3.7) $[v]_{h}(t):=e^{-\frac{t}{h}}v_{o}+\tfrac{1}{h}\int_{0}^{t}e^{\frac{s-t}{h}}v(s)\,\mathrm{d}s$
for any $t\in[0,T]$. The basic properties of this mollification procedure are collected in the following lemma. For the proofs of the statements, we refer to [13, Appendix B] and [46].
@@
Let $X$ be a separable Banach space and $v_{o}\in X$. If $v\in L^{r}(0,T;X)$ for some $1\leq r\leq\infty$, then the mollification $[v]_{h}$ defined according to (3.7) satisfies $[v]_{h}\in L^{r}(0,T;X)$ and for any $t_{o}\in(0,T]$ there holds the bound
@@
where the bracket $[\ldots]^{\frac{1}{r}}$ is interpreted as $1$ in the case $r=\infty$. Furthermore, in the case $r<\infty$ we have that $[v]_{h}\to v$ in $L^{r}(0,T;X)$ in the limit $h\downarrow 0$. If $v\in C^{0}([0,T];X)$ and $v_{o}=v(0)$, then $[v]_{h}\in C^{0}([0,T];X)$ with $[v]_{h}(0)=v_{o}$ and moreover, we have that $[v]_{h}\to v$ in $C^{0}([0,T];X)$ as $h\downarrow 0$.
Next, we recall the results from [16, Lemma 2.3], which yields the convergence $f(x,[v]_{h},D[v]_{h})\to f(x,v,Dv)$ as $h\downarrow 0$, provided that (2.4) holds. Note that the property (3.8) is evident from the proof of [16, Lemma 2.3].
@@ 3.4. Continuity in time
Since (2.2) holds, it follows that $\operatorname{spt}\big{(}u_{o}^{(\varepsilon)}\big{)}\subset E^{t}$ for all $t\in[0,T)$. Now, let $[u]_{\lambda,\varepsilon}$ with $\lambda>0$ be the time mollification of $u$ according to (3.7) with initial values $u_{o}^{(\varepsilon)}$. Using $[u]_{\lambda,\varepsilon}$ as a comparison map in the variational inequality  (2.8), we obtain that
@@
$\leq\iint_{E\cap\Omega_{\tau}}f\big{(}x,[u]_{\lambda,\varepsilon},D[u]_{\lambda,\varepsilon}\big{)}\,\mathrm{d}x\mathrm{d}t+\iint_{E\cap\Omega_{\tau}}\partial_{t}[u]_{\lambda,\varepsilon}\cdot\big{(}\boldsymbol{[u]_{\lambda,\varepsilon}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t$
@@
$\leq\iint_{E\cap\Omega_{\tau}}\big{[}f(x,u,Du)\big{]}_{\lambda,\varepsilon}\,\mathrm{d}x\mathrm{d}t-\tfrac{1}{h_{\ell}}\iint_{E\cap\Omega_{\tau}}\big{(}[u]_{\lambda,\varepsilon}-u\big{)}\cdot\big{(}\boldsymbol{[u]_{\lambda,\varepsilon}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\,\mathrm{d}x\mathrm{d}t$
@@
Since the right-hand side vanishes as $i\to\infty$, and since we have that $[u]_{\lambda_{i},\varepsilon_{i}}\in C^{0}([0,T];L^{q+1}(\Omega,\mathds{R}^{N}))$ for all $i\in\mathds{N}$, this implies the claim of the lemma. ∎
@@ 3.5. Density of smooth functions
Note that in the special case of nondecreasing domains, (3.19) is always satisfied. In the following, we generalize the corresponding statements for the case $q=1$ in [17, Section 5.3.1]. First, we define the space
@@
$\leq\operatorname{dist}(\widetilde{x},\Omega\setminus E^{s})+|y-\widetilde{x}|$
$\leq\boldsymbol{e^{\mbox{\unboldmath{\scriptsize$c$}}}}(E^{s},E^{t})+|y-x|+|x-\widetilde{x}|$
@@
Further, let $\widetilde{\eta}\in C^{0,1}(\mathds{R})$ be a cut-off function such that $\widetilde{\eta}\equiv 0$ in $(-\infty,1]$, $\widetilde{\eta}(r):=r-1$ for $r\in(1,2)$ and $\widetilde{\eta}\equiv 1$ in $[2,\infty)$. For $\sigma>0$, we define
@@
Applying Corollary 3.20 with the choice $q=p-1$, we deduce that $C_{0}^{\infty}(E,\mathds{R}^{N})$ is dense in $V^{p,0}(E)$ with respect to the norm topology in $V^{p,0}(E)$.
@@ 4.1. Time discretization
$\left\{\begin{array}[]{ll}u_{j}\rightharpoondown u&\text{weakly in $W^{1,p}(\Omega,\mathds{R}^{N})$ as $j\to\infty$,}\\[5.0pt] u_{j}\rightharpoondown u&\text{weakly in $L^{q+1}(\Omega,\mathds{R}^{N})$ as $j\to\infty$.}\end{array}\right.$
@@
Since $\mathcal{F}_{h}[\cdot,\overline{u}]$ is convex, by [19, Corollary 3.9] $\mathcal{F}_{h}$ is lower semicontinuous also with respect to weak convergence in $L^{q+1}(\Omega,\mathds{R}^{N})\cap W^{1,p}(\Omega,\mathds{R}^{N})$. This gives us that
@@ 4.2. Reformulation of the minimizing property
$\phantom{=}+\tfrac{1}{h_{\ell}}\int_{\Omega}\tfrac{1}{q+1}\big{(}|w_{s}|^{q+1}-|u_{\ell,i}|^{q+1}\big{)}-s\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot(w-u_{\ell,i})\,\mathrm{d}x.$
@@
$\phantom{=}+\tfrac{1}{h_{\ell}}\int_{\Omega}\tfrac{1}{(q+1)s}\big{(}|w_{s}|^{q+1}-|u_{\ell,i}|^{q+1}\big{)}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot(w-u_{\ell,i})\,\mathrm{d}x.$
Since $\frac{1}{q+1}|\cdot|^{q+1}$ is convex, the map $s\mapsto\tfrac{1}{(q+1)s}\big{(}|w_{s}|^{q+1}-|u_{\ell,i}|^{q+1}\big{)}$ is monotone and converges a.e. in $\Omega$ to $\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot(w-u_{\ell,i})$ in the limit $s\downarrow 0$. Therefore, passing to the limit in the preceding inequality by means of the dominated convergence theorem shows that
@@
(4.2) $\leq\int_{\Omega}f(x,w,Dw)\,\mathrm{d}x+\tfrac{1}{h_{\ell}}\int_{\Omega}\big{(}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot(w-u_{\ell,i})\,\mathrm{d}x$
@@ 4.3. Energy estimates and weak convergence
(4.3) $\big{|}\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\cdot u_{\ell,i}\big{|}\leq|u_{\ell,i-1}|^{q}|u_{\ell,i}|\leq\tfrac{q}{q+1}|u_{\ell,i-1}|^{q+1}+\tfrac{1}{q+1}|u_{\ell,i}|^{q+1}$
@@
$\leq h_{\ell}\int_{\Omega}f(x,u_{\ast},Du_{\ast})\,\mathrm{d}x+\tfrac{q}{q+1}\int_{\Omega}|u_{\ell,i-1}|^{q+1}\,\mathrm{d}x+\int_{\Omega}\big{(}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot u_{\ast}\,\mathrm{d}x$
@@
$\leq mh_{\ell}\int_{\Omega}f(x,u_{\ast},Du_{\ast})\,\mathrm{d}x+\tfrac{q}{q+1}\int_{\Omega}|u_{o}|^{q+1}\,\mathrm{d}x+\int_{\Omega}\big{(}\boldsymbol{u_{\ell}(mh_{\ell})^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{o}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot u_{\ast}\,\mathrm{d}x$
@@
(4.6) $\left\{\begin{array}[]{ll}u_{\ell}\overset{\raisebox{-4.30554pt}{\scriptsize$*$}}{\rightharpoondown}u&\text{weakly${}^{\ast}$ in $L^{\infty}(0,T;L^{q+1}(\Omega,\mathds{R}^{N}))$ as $\mathfrak{K}\ni\ell\to\infty$,}\\[5.0pt] u_{\ell}\rightharpoondown u&\text{weakly in $L^{p}\big{(}0,T;W^{1,p}_{u_{\ast}}(\Omega,\mathds{R}^{N})\big{)}$ as $\mathfrak{K}\ni\ell\to\infty$.}\end{array}\right.$
@@ 4.4. Boundary values
In this section, we show that $u\in V_{q}^{p}(E)$, i.e. that the limit map admits the correct boundary values. This means that $u(t)=u_{\ast}$ a.e. in $\Omega\setminus E^{t}$. We proceed as in [17, Section 4.1.4]. Let $\ell\in\mathds{N}$ and define
@@ 4.5. Convergence almost everywhere
In this section, we analyze the convergence of $u_{\ell}$ to $u$ further by proceeding as in [14, Section 6.3]. To this end, note that $w=u_{\ell,i-1}$ is an admissible comparison map in (4.2), since $u_{\ell,i-1}\in V_{ih_{\ell}}$ for $i=2,\ldots,\ell$ by (2.2). Together with Lemma 3.4, this gives us that
$\int_{\Omega}$ $f(x,u_{\ell,i},Du_{\ell,i})\,\mathrm{d}x+\tfrac{c(q)}{h_{\ell}}\int_{\Omega}\Big{|}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}^{2}\,\mathrm{d}x$
$\leq\int_{\Omega}f(x,u_{\ell,i},Du_{\ell,i})\,\mathrm{d}x+\tfrac{1}{h_{\ell}}\int_{\Omega}\big{(}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$q$}}}}\big{)}\cdot(u_{\ell,i}-u_{\ell,i-1})\,\mathrm{d}x$
@@
$\int_{\Omega}$ $f(x,u_{\ell,i_{2}},Du_{\ell,i_{2}})\,\mathrm{d}x+\tfrac{c(q)}{h_{\ell}}\sum_{i=i_{1}+1}^{i_{2}}\int_{\Omega}\Big{|}\boldsymbol{u_{\ell,i}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}^{2}\,\mathrm{d}x$
@@
$\int_{\Omega}$ $f(x,u_{\ell}(i_{2}h_{\ell}),Du_{\ell}(i_{2}h_{\ell}))\,\mathrm{d}x+c(q)\int_{i_{1}h_{\ell}}^{i_{2}h_{\ell}}\int_{\Omega}\Big{|}\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}^{2}\,\mathrm{d}x$
@@
$\int_{\Omega}$ $f(x,u_{\ell}(\tau),Du_{\ell}(\tau))\,\mathrm{d}x+c(q)\iint_{\Omega\times(\varepsilon,\tau)}\Big{|}\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\Big{|}^{2}\,\mathrm{d}x$
[…]
@@
(4.10) $\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\to\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\quad\text{strongly in $L^{1}\big{(}\Omega\times(\varepsilon,T),\mathds{R}^{N}\big{)}$ as$\mathfrak{K}\ni\ell\to\infty$}.$
@@
Moreover, we will now have a closer look on the time derivative of our constructed solutions. Omitting the first term on the left-hand side in (4.8) gives us an estimate for the finite difference quotient of $\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}$. Therefore, there exists a subsequence, still denoted by $\mathfrak{K}\subset\mathds{N}$, such that $\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\rightharpoondown w$ in $L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ as $\mathfrak{K}\ni\ell\to\infty$. Combining this with (4.10) yields for all $\phi\in C_{0}^{\infty}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ that
$\iint_{\Omega\times(\varepsilon,T)}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\cdot\partial_{t}\phi\,\mathrm{d}x\mathrm{d}t$ $=\lim_{\mathfrak{K}\ni\ell\to\infty}\iint_{\Omega\times(\varepsilon,T)}\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\cdot\Delta_{h_{\ell}}\phi\,\mathrm{d}x\mathrm{d}t$
$=-\lim_{\mathfrak{K}\ni l\to\infty}\iint_{\Omega\times(\varepsilon,T)}\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\cdot\phi\,\mathrm{d}x\mathrm{d}t$
@@
From this we conclude that $w=\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$. Furthermore, since we have that $\partial_{t}\boldsymbol{|u|^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ by the chain rule, it follows that
$\partial_{t}\boldsymbol{u}^{q+1}=\partial_{t}\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}|u|^{\frac{q+1}{2}}+\boldsymbol{u^{\mbox{\unboldmath{\scriptsize$\frac{q+1}{2}$}}}}\partial_{t}|u|^{\frac{q+1}{2}}\in L^{1}\big{(}\Omega\times(\varepsilon,T),\mathds{R}^{N}\big{)}.$

### v2

@@ 3.1. Technical lemmas
(3.5) $\left\{\begin{array}[]{ll}\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\to\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}&\text{strongly in }L^{1}(\Omega_{T},\mathds{R}^{N}),\\[5.0pt] u_{\ell}\to u&\text{a.e.~in }\Omega_{T},\end{array}\right$.
@@ 3.2. Mollification in time
We will use the following regularization with respect to time introduced by Landes [41]. Given $h>0$ and a function $v$, the mollification $v]_{h}$ is constructed to satisfy the ordinary differential equation
@@
with initial condition $v]_{h}(0)=v_{o}$. More precisely, let $X$ be a separable Banach space, $v_{o}\in X$ and $v\in L^{r}(0,T;X)$ for some $1\leq r\leq\infty$. Later on, we will mainly consider $X=L^{q}(\Omega,\mathds{R}^{N})$ and $X=W^{1,p}_{u_{\ast}}(\Omega,\mathds{R}^{N})$. For $h\in(0,T]$, we define the mollification
(3.7) $v]_{h}(t):=e^{-\frac{t}{h}}v_{o}+\tfrac{1}{h}\int_{0}^{t}e^{\frac{s-t}{h}}v(s)\,\mathrm{d}s$
for any $t\in[0,T]$. The basic properties of this mollification procedure are collected in the following lemma. For the proofs of the statements, we refer to [15, Appendix B] and [48].
@@
Let $X$ be a separable Banach space and $v_{o}\in X$. If $v\in L^{r}(0,T;X)$ for some $1\leq r\leq\infty$, then the mollification $v]_{h}$ defined according to (3.7) satisfies $v]_{h}\in L^{r}(0,T;X)$ and for any $t_{o}\in(0,T]$ there holds the bound
@@
where the bracket $\ldots]^{\frac{1}{r}}$ is interpreted as $1$ in the case $r=\infty$. Furthermore, in the case $r<\infty$ we have that $v]_{h}\to v$ in $L^{r}(0,T;X)$ in the limit $h\downarrow 0$. If $v\in C^{0}([0,T];X)$ and $v_{o}=v(0)$, then $v]_{h}\in C^{0}([0,T];X)$ with $v]_{h}(0)=v_{o}$ and moreover, we have that $v]_{h}\to v$ in $C^{0}([0,T];X)$ as $h\downarrow 0$.
Next, we recall the results from [18, Lemma 2.3], which yields the convergence $f(x,[v]_{h},D[v]_{h})\to f(x,v,Dv)$ as $h\downarrow 0$, provided that (2.4) holds. Note that the property (3.8) is evident from the proof of [18, Lemma 2.3].
@@ 3.4. Continuity in time
Since (2.2) holds, it follows that $\operatorname{spt}\big(u_{o}^{(\varepsilon)}\big)\subset E^{t}$ for all $t\in[0,T)$. Now, let $u]_{\lambda,\varepsilon}$ with $\lambda>0$ be the time mollification of $u$ according to (3.7) with initial values $u_{o}^{(\varepsilon)}$. Using $u]_{\lambda,\varepsilon}$ as a comparison map in the variational inequality  (2.8), we obtain that
@@
$\leq\iint_{E\cap\Omega_{\tau}}f\big(x,[u]_{\lambda,\varepsilon},D[u]_{\lambda,\varepsilon}\big)\,\mathrm{d}x\mathrm{d}t+\iint_{E\cap\Omega_{\tau}}\partial_{t}[u]_{\lambda,\varepsilon}\cdot\big(\boldsymbol{[u]_{\lambda,\varepsilon}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t$
@@
$\leq\iint_{E\cap\Omega_{\tau}}\big[f(x,u,Du)\big]_{\lambda,\varepsilon}\,\mathrm{d}x\mathrm{d}t-\tfrac{1}{h_{\ell}}\iint_{E\cap\Omega_{\tau}}\big([u]_{\lambda,\varepsilon}-u\big)\cdot\big(\boldsymbol{[u]_{\lambda,\varepsilon}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u^{\mbox{{\scriptsize$q$}}}}\big)\,\mathrm{d}x\mathrm{d}t$
@@
Since the right-hand side vanishes as $i\to\infty$, and since we have that $u]_{\lambda_{i},\varepsilon_{i}}\in C^{0}([0,T];L^{q+1}(\Omega,\mathds{R}^{N}))$ for all $i\in\mathds{N}$, this implies the claim of the lemma. ∎
@@ 3.5. Density of smooth functions
Note that in the special case of nondecreasing domains, (3.19) is always satisfied. In the following, we generalize the corresponding statements for the case $q=1$ in [19, Section 5.3.1]. First, we define the space
@@
[…]
@@
Further, let $\widetilde{\eta}\in C^{0,1}(\mathds{R})$ be a cut-off function such that $\widetilde{\eta}\equiv 0$ in $(-\infty,1]$, $\widetilde{\eta}(r):=r-1$ for $r\in(1,2)$ and $\widetilde{\eta}\equiv 1$ in $2,\infty)$. For $\sigma>0$, we define
@@
Applying Corollary 3.20 with the choice q=p1q=p-1, we deduce that $C_{0}^{\infty}(E,\mathds{R}^{N})$ is dense in $V^{p,0}(E)$ with respect to the norm topology in $V^{p,0}(E)$.
@@ 4.1. Time discretization
[…]
@@
Since $\mathcal{F}_{h}[\cdot,\overline{u}]$ is convex, by [21, Corollary 3.9] $\mathcal{F}_{h}$ is lower semicontinuous also with respect to weak convergence in $L^{q+1}(\Omega,\mathds{R}^{N})\cap W^{1,p}(\Omega,\mathds{R}^{N})$. This gives us that
@@ 4.2. Reformulation of the minimizing property
$\phantom{=}+\tfrac{1}{h_{\ell}}\int_{\Omega}\tfrac{1}{q+1}\big(|w_{s}|^{q+1}-|u_{\ell,i}|^{q+1}\big)-s\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\cdot(w-u_{\ell,i})\,\mathrm{d}x.$
@@
$\phantom{=}+\tfrac{1}{h_{\ell}}\int_{\Omega}\tfrac{1}{(q+1)s}\big(|w_{s}|^{q+1}-|u_{\ell,i}|^{q+1}\big)-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\cdot(w-u_{\ell,i})\,\mathrm{d}x.$
Since $\frac{1}{q+1}|\cdot|^{q+1}$ is convex, the map $s\mapsto\tfrac{1}{(q+1)s}\big(|w_{s}|^{q+1}-|u_{\ell,i}|^{q+1}\big)$ is monotone and converges a.e. in $\Omega$ to $\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$q$}}}}\cdot(w-u_{\ell,i})$ in the limit $s\downarrow 0$. Therefore, passing to the limit in the preceding inequality by means of the dominated convergence theorem shows that
@@
(4.2) $\leq\int_{\Omega}f(x,w,Dw)\,\mathrm{d}x+\tfrac{1}{h_{\ell}}\int_{\Omega}\big(\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\big)\cdot(w-u_{\ell,i})\,\mathrm{d}x$
@@ 4.3. Energy estimates and weak convergence
(4.3) $\big|\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\cdot u_{\ell,i}\big|\leq|u_{\ell,i-1}|^{q}|u_{\ell,i}|\leq\tfrac{q}{q+1}|u_{\ell,i-1}|^{q+1}+\tfrac{1}{q+1}|u_{\ell,i}|^{q+1}$
@@
$\leq h_{\ell}\int_{\Omega}f(x,u_{\ast},Du_{\ast})\,\mathrm{d}x+\tfrac{q}{q+1}\int_{\Omega}|u_{\ell,i-1}|^{q+1}\,\mathrm{d}x+\int_{\Omega}\big(\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\big)\cdot u_{\ast}\,\mathrm{d}x$
@@
$\leq mh_{\ell}\int_{\Omega}f(x,u_{\ast},Du_{\ast})\,\mathrm{d}x+\tfrac{q}{q+1}\int_{\Omega}|u_{o}|^{q+1}\,\mathrm{d}x+\int_{\Omega}\big(\boldsymbol{u_{\ell}(mh_{\ell})^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{o}^{\mbox{{\scriptsize$q$}}}}\big)\cdot u_{\ast}\,\mathrm{d}x$
@@
[…]
@@ 4.4. Boundary values
In this section, we show that $u\in V_{q}^{p}(E)$, i.e. that the limit map admits the correct boundary values. This means that $u(t)=u_{\ast}$ a.e. in $\Omega\setminus E^{t}$. We proceed as in [19, Section 4.1.4]. Let $\ell\in\mathds{N}$ and define
@@ 4.5. Convergence almost everywhere
In this section, we analyze the convergence of $u_{\ell}$ to $u$ further by proceeding as in [16, Section 6.3]. To this end, note that $w=u_{\ell,i-1}$ is an admissible comparison map in (4.2), since $u_{\ell,i-1}\in V_{ih_{\ell}}$ for $i=2,\ldots,\ell$ by (2.2). Together with Lemma 3.4, this gives us that
$\int_{\Omega}$ $f(x,u_{\ell,i},Du_{\ell,i})\,\mathrm{d}x+\tfrac{c(q)}{h_{\ell}}\int_{\Omega}\Big|\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|^{2}\,\mathrm{d}x$
$\leq\int_{\Omega}f(x,u_{\ell,i},Du_{\ell,i})\,\mathrm{d}x+\tfrac{1}{h_{\ell}}\int_{\Omega}\big(\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$q$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$q$}}}}\big)\cdot(u_{\ell,i}-u_{\ell,i-1})\,\mathrm{d}x$
@@
$\int_{\Omega}$ $f(x,u_{\ell,i_{2}},Du_{\ell,i_{2}})\,\mathrm{d}x+\tfrac{c(q)}{h_{\ell}}\sum_{i=i_{1}+1}^{i_{2}}\int_{\Omega}\Big|\boldsymbol{u_{\ell,i}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}-\boldsymbol{u_{\ell,i-1}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|^{2}\,\mathrm{d}x$
@@
$\int_{\Omega}$ $f(x,u_{\ell}(i_{2}h_{\ell}),Du_{\ell}(i_{2}h_{\ell}))\,\mathrm{d}x+c(q)\int_{i_{1}h_{\ell}}^{i_{2}h_{\ell}}\int_{\Omega}\Big|\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|^{2}\,\mathrm{d}x$
@@
$\int_{\Omega}$ $f(x,u_{\ell}(\tau),Du_{\ell}(\tau))\,\mathrm{d}x+c(q)\iint_{\Omega\times(\varepsilon,\tau)}\Big|\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\Big|^{2}\,\mathrm{d}x$
$\leq\mathchoice{\vbox{\hbox{$\textstyle-$}}\kern-4.86108pt}{\vbox{\hbox{$\scriptstyle-$}}\kern-3.43057pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.908pt}{\vbox{\hbox{$\scriptscriptstyle-$}}\kern-2.76045pt}\!\int_{h_{\ell}}^{\varepsilon-h_{\ell}}\int_{\Omega}f(x,u_{\ell},Du_{\ell})\,\mathrm{d}x\mathrm{d}t$
@@
(4.10) $\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\to\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\quad\text{strongly in $L^{1}\big(\Omega\times(\varepsilon,T),\mathds{R}^{N}\big)$ as $\mathfrak{K}\ni\ell\to\infty$}.$
@@
Moreover, we will now have a closer look on the time derivative of our constructed solutions. Omitting the first term on the left-hand side in (4.8) gives us an estimate for the finite difference quotient of $\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}$. Therefore, there exists a subsequence, still denoted by $\mathfrak{K}\subset\mathds{N}$, such that $\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\rightharpoondown w$ in $L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ as $\mathfrak{K}\ni\ell\to\infty$. Combining this with (4.10) yields for all $\phi\in C_{0}^{\infty}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ that
$\iint_{\Omega\times(\varepsilon,T)}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\cdot\partial_{t}\phi\,\mathrm{d}x\mathrm{d}t$ $=\lim_{\mathfrak{K}\ni\ell\to\infty}\iint_{\Omega\times(\varepsilon,T)}\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\cdot\Delta_{h_{\ell}}\phi\,\mathrm{d}x\mathrm{d}t$
$=-\lim_{\mathfrak{K}\ni l\to\infty}\iint_{\Omega\times(\varepsilon,T)}\Delta_{-h_{\ell}}\boldsymbol{u_{\ell}^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\cdot\phi\,\mathrm{d}x\mathrm{d}t$
@@
From this we conclude that $w=\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$. Furthermore, since we have that $\partial_{t}\boldsymbol{|u|^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\in L^{2}(\Omega\times(\varepsilon,T),\mathds{R}^{N})$ by the chain rule, it follows that
$\partial_{t}\boldsymbol{u}^{q+1}=\partial_{t}\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}|u|^{\frac{q+1}{2}}+\boldsymbol{u^{\mbox{{\scriptsize$\frac{q+1}{2}$}}}}\partial_{t}|u|^{\frac{q+1}{2}}\in L^{1}\big(\Omega\times(\varepsilon,T),\mathds{R}^{N}\big).$